How many square yards of linoleum are needed for a kitchen floor that measures 10 ft by 12 ft?
A sheet of plastic 120 cm by 76 cm is to be cut for desks tops that measure 60 cm long and 38 cm wide. How many tops can be made from 6 sheets?
Find the area of a square whose perimeter is 7.2 m.
What is the area of a rectangle 8 m long if its perimeter is 30 m?
Thanks!!!

Answer by feliz1965(151)

What about placing one question per post?
By doing so, it gives many tutors the chance to reply to your questions.
There is more than one way to answer math questions. So, by giving many tutors the chance to reply, you may find a method that is to your liking.
I will do 2 questions only.
(1) How many square yards of linoleum are needed for a kitchen floor that measures 10 ft by 12 ft?
To answer this question, we need to convert feet to yards.
10 feet times 12 feet = 120 feet.
How many yards in 120 feet?
There are 3 feet in one yard.
So, 120 feet divided by 3 feet = 40.
The answer is 40 yards.
==================================================

(2) Find the area of a square whose perimeter is 7.2 m.
A square has equal measure on each side.
A square has 4 sides.
We are given the perimeter to be 7.2 meters.
To find one side, we divide 7.2 by 4.
Side of this square = 7.2/4
Side of this square = 1.8 meters.
We now use the formula for the area of a square, which is Area = side squared or in short A = s^2.
We plug 1.8 for s, square it and there is your area.
A = (1.8)^2
A = 1.8 times 1.8
A = 3.24 meters^2
The little 2 after the word "meters" indicates that I found the area because area is always measured in square units.